Show the energy consideration of harmonic oscillator at any time t and show that the energy decreases exponentially at a faster rate , E(t)= E0e-2βt  where E0= 1/2KA2

1
Expert's answer
2021-01-11T11:42:47-0500

Answer

harmonic oscillator differential equation is given by

"\frac{md^2x}{dt^2}+kx=0"

Solution of above euation can be calculated by differential equation

Which is found

"x=Ae^{-\beta t}"

Differentiate with respect to t

"v=\frac{dx}{dt}=-A\beta e ^{-\beta t}"

the energy consideration of harmonic oscillator at any time t is given by

"E=\frac{mv^2}{2}"

Putting value of velocity

"E=\frac{e^{-2\beta t}}{2KA^2}=E_0e^{-2\beta t}"

Where

"E_0= 1\/2KA^2"